Circular No. 8491

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVAE 2005am AND 2005an Further to IAUC 8484, K. Shimasaki and W. Li report the LOSS dis-

covery of an apparent supernova on an unfiltered KAIT image taken on Mar. 3.35 UT (mag 17.3). SN 2005an is located at α = 12^{h}28^{m}27^{s}.35, δ = −24^{o}38′02′′.4 (equinox 2000.0), which is 0′′.2 east and 10′′.3 north of the nucleus of ESO 506-G11. A KAIT image taken on Jan. 20.36 showed nothing at this position (limiting mag 19.0).

M. Modjaz, R. Kirshner and P. Challis, Harvard-Smithsonian Center for Astrophysics, report that a spectrum (range 350–740 nm) of SN 2005an, obtained on Mar. 4.38 UT by H. Hao with the F. L. Whipple Observatory 1.5-m telescope (+ FAST), shows it to be most probably a young type-II supernova. The noisy spectrum consists of a featureless and blue continuum and is similar to an early spectrum of SN 1993J (Matheson et al. 2000, A.J. 120, 1487). A spectrum of SN 2005am (cf. IAUC 8490), obtained on Mar. 3.29, shows it to be a type-Ia supernova, one or two weeks before maximum. The supernova expansion velocity, derived from the minimum of Si II (rest 635.5 nm) and adopting the NED recession velocity of 2368 km/s for the host galaxy, is ∼ 12700 km/s.

COMET C/2005 E1 (TUBBIOLO) A. F. Tubbiolo reports his discovery of a comet on images taken with

the 0.9-m Spacewatch reflector at Kitt Peak (discovery observation given below), noting a tail ≈ 13′′ long to the northwest on the discovery images, with a tail and coma noted on CCD images taken on Mar. 4.41–4.45 UT. Following posting on the ‘NEO Confirmation Page’, co-added R-band images taken by A. Fitzsimmons, S. Lowry, and C. Snodgrass with the 2-m ‘Faulkes Telescope North’ at Haleakala on Mar. 4.4 show a faint tail 15′′ long in p.a. 300^{o}. CCD images taken by M. Tichý and J. Tichá with the 1.06-m telescope at Kleť on Mar. 4.9 show the comet to be diffuse.

2005 UT  α_{2000}  δ_{2000}  Mag.
Mar. 3.25502  11^{h}48^{m}54.^{s}36  +2^{o}34′40′′.5  20.9

The available astrometry, the following indeterminate parabolic orbital elements, an ephemeris appear on MPEC 2005-E35.

T = 2006 Nov. 15.867 TT  ω = 296.^{o}728
Ω = 357.877 }2000.0
q = 1.66913 AU  i = 17.439
2005 March 7  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2005 CBAT  Daniel W. E. Green